Home‎ > ‎Crew 161‎ > ‎

0111-DailySummaryReport

MDRS Daily Summary Report

Sol 2 1/11/2016

First EVA!

Executive Officer Adeniyi Adebayo Reporting

Mission Status: Today, the crew got off to our first-ever EVA. Soil samples were successfully collected and plant studies have started in earnest.

Sol Activity Summary: The highlight of the day was EVA #1. Three crew members headed north of the Hab to collect soil samples of the red planet. Our biologist has had her hands full. There was a minor glitch with the GPS on the EVA team, but HabComm was on the radio and guided the team back to safety. Team debrief followed as soon as the EVA crew returned and each crew member’s role was evaluated with a view to identifying ways to improve performance on future EVAs.

Look Ahead Plan: An EVA is planned for tomorrow. A team of three will go out for more sample collections and mapping out territories around the hab.

Anomalies in work: There are no anomalies to report.

Weather: The Martian sky was very bright and clear today. Our EVA crew reported a warm weather.

Crew Physical Status: All crew members are doing fine.

EVA: Today’s EVA was for soil sample collection. This was successfully done and the EVA team returned safely back to the Hab.

Reports to be filed:
Commanders Report
Daily Summary Report
Journalist Report
Biologist Report
Engineering Report

Support Requested: Approval for EVA for SOL 3